May someone please help me out with this
EQUATION 9.2:
C = 400 + 0.6Yd
Taxes = 600
Equilibrium Output = $4,000
39) Refer to Equation 9.2. At equilibrium, the sum of investment and government purchases in Peru is
40) Refer to Equation 9.2. At equilibrium, government purchases in Peru is
41) Refer to Equation 9.2. At equilibrium, saving in Peru equals
42) Refer to Equation 9.2. At equilibrium, leakages in Peru equal
1
Expert's answer
2020-04-02T09:52:44-0400

Yd=Y-T=4 000-600=3 400

C=400+0.6xYd=400+0.6x3400=400+2040=2440

39) Y=C+I+G

4000=2440+I+G

1560=I+G

40) I=S

S=-400+0.6xYd=-400+2040=1640

G=-80, 1560=I+G

41) C+I>Y

2440+1640>4000

4080>400, 80

42)leakages S=I=1640